# Studying geometry of the ultraluminous X-ray pulsar Swift J0243.6+6124 using X-ray and optical polarimetry

### v1

@@
Discovery of pulsations from a number of ultra-luminous X-ray (ULX) sources proved that accretion onto neutron stars can produce luminosities exceeding the Eddington limit by a couple of orders of magnitude. The conditions necessary to achieve such high luminosities as well as the exact geometry of the accretion flow in the neutron star vicinity are, however, a matter of debate. The pulse phase-resolved polarization measurements that became possible with the launch of the Imaging X-ray Polarimetry Explorer (IXPE) can be used to determine the pulsar geometry and its orientation relative to the orbital plane. They provide an avenue to test different theoretical models of ULX pulsars. In this paper we present the results of three IXPE observations of the first Galactic ULX pulsar Swift J0243.6+6124 during its 2023 outburst. We find strong variations of the polarization characteristics with the pulsar phase. The average polarization degree (PD) increases from about 5% to 15% as the flux dropped by a factor of three in the course of the outburst. The polarization angle (PA) as function of the pulsar phase shows two peaks in the first two observations, but changes to a characteristic sawtooth pattern in the remaining data set. This is not consistent with a simple rotating vector model. Assuming the existence of an additional constant polarized component, we were able to fit the three observations with a common rotating vector model and obtain constraints on the pulsar geometry. In particular, we find the pulsar angular momentum inclination with respect to the line-of-sight of 15°– 40°, the magnetic obliquity of 60°– 80°, and the pulsar spin position angle of , which significantly differs from the constant component PA of about 10°. Combining these X-ray measurements with the optical PA, we find evidence for at least a 30° misalignment between the pulsar angular momentum and the binary orbital axis.
@@
The accretion flow in the vicinity of the NS surface is governed by the magnetic field which channels the accreting matter towards the magnetic poles. At low accretion rates, the gas decelerates at the NS surface forming hotspots that radiate in the X-ray band (Zel’dovich & Shakura, 1969). At high accretion rates, the radiation-dominated shock stops the gas above the NS surface forming accretion columns (Basko & Sunyaev, 1976; Lyubarskii & Syunyaev, 1988; Mushtukov et al., 2015). The ULX pulsars are characterized by luminosities exceeding the Eddington limit by a factor of 10–500. It is possible to achieve such a high luminosity thanks to a reduction in the interaction cross-section between matter and radiation in a strong magnetic field (Mushtukov et al., 2015), possibly dominated by quadrupole component (Tsygankov et al., 2018; Brice et al., 2021). Alternatively, a strong beaming of radiation by disk outflows can be responsible for high apparent luminosities similarly to ULXs hosting black holes (King et al., 2001; Poutanen et al., 2007). However, the latter scenario meet obstacles in the case of strongly magnetized NSs, because the size of the magnetosphere exceeds the NS size by two orders of magnitude, reducing the strength of the outflow (Lipunov, 1982; Chashkina et al., 2019). Furthermore, the observed radiation cannot be strongly beamed because of the large pulsation amplitude (Mushtukov et al., 2021; Mushtukov & Portegies Zwart, 2023), implying enormous accretion rates g s-1.
@@
Complexities in ULX pulsars studies arise from their considerable distances in external galaxies. Discovery of the first Galactic ULX pulsar Swift J0243.6+6124 (J0243 hereafter) in 2017 opened new exciting possibilities. J0243 was detected by the Swift/BAT (Cenko et al., 2017) and soon after pulsations at a period of 9.86 s were found with the Swift/XRT (Kennea et al., 2017). Kong et al. (2022) reported discovery of a cyclotron line scattering feature at the energy 130 keV, which implies strong surface magnetic field of G. An optical counterpart, a late Oe-type or early Be-type star, was identified as USNO-B1.0 1514+0083050 based on positional coincidence with the Swift/XRT source (Kennea et al., 2017; Bikmaev et al., 2017; Kouroubatzakis et al., 2017). This star appears in the Gaia Data Release 3 (DR3) catalog (id: 465628193526364416) at the distance determined via parallax of 5.2 kpc111We note that this is closer than the distance of 6.8 kpc from the Gaia DR2 catalog (Bailer-Jones et al., 2018) and used in some earlier literature. (Bailer-Jones et al., 2021). Using this distance, the peak luminosity in the 2017 outburst was erg s-1 (Doroshenko et al., 2018; Tsygankov et al., 2018; Wilson-Hodge et al., 2018).
@@
Pulsar radiation was predicted to be strongly polarized (Meszaros et al., 1988). Testing this prediction became possible recently thanks to the Imaging X-ray Polarimetry Explorer (IXPE) that allows us to measure X-ray polarization in the 2–8 keV energy band. IXPE discovered pulse-phase dependent polarization in a number of X-ray pulsars (e.g., Doroshenko et al., 2022, 2023; Tsygankov et al., 2022, 2023; Forsblom et al., 2023; Malacaria et al., 2023; Suleimanov et al., 2023; Mushtukov et al., 2023; Heyl et al., 2024). Variations of the polarization properties with the pulsar phase allow us to constrain the pulsar geometry.
@@
During the peak of the outburst, spectrum of J0243 was dominated by the Compton reflection component with a strong iron line associated with reflection of the primary pulsar radiation from the well formed by the inner edge of the geometrically thick super-Eddington accretion disk (Bykov et al., 2022). This interpretation was later supported by the detection of the pulsations in the iron line with the Insight-HXMT (Xiao et al., 2024). The reflection component is also expected to be strongly polarized (Matt, 1993; Poutanen et al., 1996), providing information about the reflector geometry and observer inclination. Indeed, a strongly polarized X-ray continuum attributed to reflection was observed by IXPE in other types of compact X-ray sources: in two Seyfert 2 galaxies, the Circinus galaxy (Ursini et al., 2023) and NGC 1068 (Marin et al., 2024), as well as in the X-ray binary Cyg X-3 (Veledina et al., 2024). In all these cases, reflection is likely coming from a torus-like structure blocking the direct view towards the central X-ray source.
@@
In 2023, J0243 underwent another outburst detected first with MAXI on 2023 April 8 (Setoguchi et al., 2023). A few days later, the source was localized to be J0243 with the Swift/XRT (Kennea et al., 2023) and subsequently 9.8 s pulsations were detected with NICER (Ng et al., 2023). The source continued to brighten, reaching peak flux in early July. The light curves of J0243 as seen by the MAXI222http://maxi.riken.jp/ (Matsuoka et al., 2009) in the 2–20 keV energy band and Swift/BAT333https://swift.gsfc.nasa.gov/results/transients/ (Gehrels et al., 2004) in the 15–50 keV band during the 2023 outburst are shown in Fig. 1.
@@ 2.1 IXPE observations
IXPE is a NASA mission in partnership with the Italian space agency (see a detailed description in Weisskopf et al., 2022), launched by a Falcon 9 rocket on 2021 December 9. There are three grazing incidence telescopes onboard the observatory. Each telescope comprises an X-ray mirror assembly and a polarization-sensitive detector unit (DU) equipped with a gas-pixel detector (GPD; Soffitta et al., 2021; Baldini et al., 2021). These instruments provide imaging polarimetry in the 2–8 keV energy band with a time resolution better than 10 s.
IXPE observed J0243 during July 20–August 25 (ObsID 02250799), for three times on 2023 July 20–22, August 10–11, and August 23–25 with the total exposures of 168, 77, and 131 ks per telescope (see Fig. 1). We refer to these observations as Obs. 1, 2, and 3, respectively. The data have been processed with the ixpeobssim package version 31.0.1 (Baldini et al., 2022) using the CalDB version 20230702:v13. Source photons were collected in a circular region with a radius centered at the J0243 position. Because of the source brightness, the background was not subtracted following the recommendation by Di Marco et al. (2023). We performed the unweighted analysis (i.e. all events are taken into account independently of the quality of track reconstruction; Di Marco et al. 2022) of the IXPE data.
@@ 2.2 Timing
The event arrival times were corrected to the Solar system barycenter using the standard barycorr tool from the ftools666http://heasarc.gsfc.nasa.gov/ftools package (Blackburn, 1995) and accounting for the effects of binary motion using the orbital parameters from the Fermi/GBM777https://gammaray.nsstc.nasa.gov/gbm/science/pulsars.html (see Table 1). The spin period (of s) and the spin period derivative were then determined for each IXPE observation segment using epoch folding search, and then refined using phase-connection technique, which allows one to unambiguously define phase of each event within a given segment. The results are presented in Table 4. The gaps between individual observation segments coupled to complex evolution of intrinsic spin frequency and pulse profiles together with remaining uncertainties in orbital parameters preclude, however, direct connection of phases between the segments.
The absolute phase alignment between the segments was, therefore, determined independently. Inspection of the pulse profile shape observed by IXPE (see bottom panel of Fig. 2) reveals some common features, i.e. the minimum around phase zero, several sub-peaks within the main peak, and a peak around phase 0.9 appearing late in the outburst. The observed phases of these features can thus be used to determine absolute phase offset for each segment such that all peaks/dips (determined through fitting of gaussian functions) appear at the same pulse phase. The residual scatter can then be used to assess remaining uncertainty in final phase alignment, which we estimate to be below 1%. The phase alignment can also be checked through the comparison of the observed IXPE pulse profiles with hard X-ray pulse profiles observed by Fermi/GBM.
To do so, we used the enhanced products provided by GBM team containing Fourier coefficients of the pulse profiles for a set of time intervals with typical duration of 3 d. These are expected to change smoothly and since Fermi data contains no data gaps, the individual pulse profiles can be aligned through the cross-correlation of subsequent time intervals. The resulting pulse profile evolution is presented in the top panel of Fig. 2. We see that changes similar to those revealed by IXPE also occur in the GBM data, i.e. the secondary peak around phase 0.9 appears to be present only at lower luminosities and disappears at higher luminosities, i.e. our initial alignment using IXPE data alone proves to be robust.
@@ 2.3 Spectral analysis
For the spectral analysis, Stokes spectra for Obs. 1, 2, and 3 were extracted using the xpbin tool’s PHA1 algorithm in ixpeobbsim, giving a set of three spectra (one for each DU) per observation. The energy spectra were binned to have at least 30 counts per energy channel. The spectra corresponding to the individual observations were fitted separately with the xspec package version 12.14.0 (Arnaud, 1996) using statistics. The uncertainties are given at the 68.3% confidence level. Considering the energy resolution and energy range of IXPE, we used a simplified model consisting of an absorbed power law plus a blackbody component to fit the J0243 spectra. To account for the interstellar absorption, a multiplicative component tbabs with the abundances from Wilms et al. (2000) was applied. A re-normalization constant const was introduced to account for possible discrepancies between effective areas of the different DUs and was fixed at unity for DU1. The full spectral model thus is
The results of the spectral fitting are shown in Fig. 3 and the best-fit parameters are given in Table 5. We see that the fit quality is not good, which is caused partially by systematic errors in the instrument response and partially by the presence of a weak iron line around 6.5 keV. This does not preclude us, however, from measuring the flux in the IXPE band. For calculations of the total luminosity we estimated the bolometric correction factor for our IXPE observations from the NuSTAR spectra at similar luminosities during the 2017 outburst (Tsygankov et al., 2018; Bykov et al., 2022). The maximum bolometric luminosity of the 2023 outburst is erg s-1, a factor of 20 lower than the peak of the 2017 outburst (see penultimate row of Table 5).
@@ 2.4 X-ray polarimetric analysis
The polarimetric parameters of J0243 were extracted utilizing the pcube algorithm (xpbin tool) in the ixpeobssim package (Baldini et al., 2022), which follows the description of Kislat et al. (2015). We derive the normalized Stokes parameters and , and subsequently the polarization degree PD= and polarization angle PA=. The uncertainties are given at the 68.3% confidence level unless stated otherwise.
The data were divided into 24 separate pulse-phase bins. Polarization was found to be small in the 2–3 keV band. This motivated us to use for the following analysis only the 3–8 keV data to reduce the noise. The evolution of the normalized Stokes parameters is shown in Fig. 4 and their pulse-phase dependence is displayed in Fig. 5b,c. The PD (shown in Fig. 5d) grows with time, from less than 8% in Obs. 1, to 12% in Obs. 2 and reaches a maximum of 24% in Obs. 3. The PA (Fig. 5e) has a double peak profile in all three observations. In Obs. 1 and 2 the amplitude of variations is about 100°,888This is in strong disagreement with the recent analysis of Obs. 1 by Majumder et al. (2024), who used only 5 phase bins. while in Obs. 3 the data are consistent with two full revolution by 180° during the spin period.
@@ 2.5 Optical polarization studies
IXPE observations have been accompanied by optical polarimetric observations in the -band with the Robopol polarimeter located in the focal plane of the 1.3 m telescope of the Skinakas observatory (Greece). The observations were performed between 2023 July 28 and August 28 with multiple pointings. The data are presented in Table 9. We determined the intrinsic source polarization by subtracting the interstellar polarization using stars #1 and #2 (see Fig. 6), which are located 3′–4′ away from J0243 at a compatible distance of 5.7 kpc, as given by the Gaia EDR3 data (Bailer-Jones et al., 2021).
We also analyzed optical polarimetric measurements of the source during its previous outburst in 2017 obtained with DIPol-2 (Piirola et al., 2014) at the T60 telescope at Haleakala, Hawaii (see Table 12 in Appendix A). On average 20 measurements of polarization were taken each night. We obtained the nightly-averaged values using the -clipping method (Kosenkov et al., 2017; Kosenkov, 2021). To estimate the interstellar polarization, we observed another field star #3 (see Fig. 6) which was reasonably bright and also has a parallax similar to that of the source.

### v2

@@
Discovery of pulsations from a number of ultra-luminous X-ray (ULX) sources proved that accretion onto neutron stars can produce luminosities exceeding the Eddington limit by several orders of magnitude. The conditions necessary to achieve such high luminosities as well as the exact geometry of the accretion flow in the neutron star vicinity are, however, a matter of debate. The pulse phase-resolved polarization measurements that became possible with the launch of the Imaging X-ray Polarimetry Explorer (IXPE) can be used to determine the pulsar geometry and its orientation relative to the orbital plane. They provide an avenue to test different theoretical models of ULX pulsars. In this paper we present the results of three IXPE observations of the first Galactic ULX pulsar Swift J0243.6+6124 during its 2023 outburst. We find strong variations in the polarization characteristics with the pulsar phase. The average polarization degree increases from about 5% to 15% as the flux dropped by a factor of three in the course of the outburst. The polarization angle (PA) as a function of the pulsar phase shows two peaks in the first two observations, but changes to a characteristic sawtooth pattern in the remaining data set. This is not consistent with a simple rotating vector model. Assuming the existence of an additional constant polarized component, we were able to fit the three observations with a common rotating vector model and obtain constraints on the pulsar geometry. In particular, we find the pulsar angular momentum inclination with respect to the line of sight of 15°– 40°, the magnetic obliquity of 60°– 80°, and the pulsar spin position angle of , which significantly differs from the constant component PA of about 10°. Combining these X-ray measurements with the optical PA, we find evidence for at least a 30° misalignment between the pulsar angular momentum and the binary orbital axis.
@@
The accretion flow in the vicinity of the NS surface is governed by the magnetic field, which channels the accreting matter toward the magnetic poles. At low accretion rates, the gas decelerates at the NS surface forming hotspots that radiate in the X-ray band (Zel’dovich & Shakura, 1969). At high accretion rates, the radiation-dominated shock stops the gas above the NS surface forming accretion columns (Basko & Sunyaev, 1976; Lyubarskii & Syunyaev, 1988; Mushtukov et al., 2015). The ULX pulsars are characterized by luminosities exceeding the Eddington limit by a factor of 10–500. It is possible to achieve such a high luminosity thanks to a reduction in the interaction cross section between matter and radiation in a strong magnetic field (Mushtukov et al., 2015), possibly dominated by quadrupole component (Tsygankov et al., 2018; Brice et al., 2021). Alternatively, a strong beaming of radiation by disk outflows can be responsible for high apparent luminosities similarly to ULXs hosting black holes (King et al., 2001; Poutanen et al., 2007). However, the latter scenario meets obstacles in the case of strongly magnetized NSs, because the size of the magnetosphere exceeds the size of the NS by two orders of magnitude, reducing the strength of the outflow (Lipunov, 1982; Chashkina et al., 2019). Furthermore, the observed radiation cannot be strongly beamed because of the large pulsation amplitude (Mushtukov et al., 2021; Mushtukov & Portegies Zwart, 2023), implying enormous accretion rates g s-1.
@@
Complexities in ULX pulsar studies arise from their considerable distances in external galaxies. Discovery of the first Galactic ULX pulsar Swift J0243.6+6124 (hereafter J0243) in 2017 opened new exciting possibilities. J0243 was detected by the Swift/BAT (Cenko et al., 2017) and soon after pulsations at a period of 9.86 s were found with the Swift/XRT (Kennea et al., 2017). Kong et al. (2022) reported the discovery of a cyclotron line scattering feature at 130 keV, which implies a strong surface magnetic field of G. An optical counterpart, a late Oe-type or early Be-type star, was identified as USNO-B1.0 1514+0083050 based on positional coincidence with the Swift/XRT source (Kennea et al., 2017; Bikmaev et al., 2017; Kouroubatzakis et al., 2017). This star appears in the Gaia Data Release 3 (DR3) catalog (ID: 465628193526364416) at the distance determined via parallax of 5.2 kpc111This is closer than the distance of 6.8 kpc from the Gaia DR2 catalog (Bailer-Jones et al., 2018) and used in some earlier literature. (Bailer-Jones et al., 2021). Using this distance, the peak luminosity in the 2017 outburst was erg s-1 (Doroshenko et al., 2018; Tsygankov et al., 2018; Wilson-Hodge et al., 2018).
@@
Pulsar radiation was predicted to be strongly polarized (Meszaros et al., 1988). Testing this prediction recently became possible thanks to the Imaging X-ray Polarimetry Explorer (IXPE) that allows us to measure X-ray polarization in the 2–8 keV energy band. IXPE discovered pulse-phase dependent polarization in a number of X-ray pulsars (e.g., Doroshenko et al., 2022, 2023; Tsygankov et al., 2022, 2023; Forsblom et al., 2023, 2024; Malacaria et al., 2023; Suleimanov et al., 2023; Mushtukov et al., 2023; Heyl et al., 2024, see Poutanen et al. 2024 for a review). Variations in the polarization properties with the pulsar phase allowed us to constrain the pulsar geometry.
@@
During the peak of the outburst, the spectrum of J0243 was dominated by the Compton reflection component with a strong iron line associated with the reflection of the primary pulsar radiation from the well formed by the inner edge of the geometrically thick super-Eddington accretion disk (Bykov et al., 2022). This interpretation was later supported by the detection of the pulsations in the iron line with the Insight-HXMT (Xiao et al., 2024). The reflection component is also expected to be strongly polarized (Matt, 1993; Poutanen et al., 1996), providing information about the reflector geometry and its orientation with respect to the observer. A strongly polarized X-ray continuum attributed to reflection was observed by IXPE in other types of compact X-ray sources: in two Seyfert 2 galaxies (the Circinus galaxy, Ursini et al. 2023; NGC 1068, Marin et al. 2024) and in the X-ray binary Cyg X-3 (Veledina et al., 2024). In all these cases the reflection is likely coming from a torus-like structure blocking the direct view toward the central X-ray source.
@@
In 2023, J0243 underwent another outburst detected first with MAXI on 2023 April 8 (Setoguchi et al., 2023). A few days later the source was localized to be J0243 with the Swift/XRT (Kennea et al., 2023), and subsequently 9.8 s pulsations were detected with NICER (Ng et al., 2023). The source continued to brighten, reaching peak flux in early July. The light curves of J0243 as seen by the MAXI222http://maxi.riken.jp/ (Matsuoka et al., 2009) in the 2–20 keV energy band and Swift/BAT333https://swift.gsfc.nasa.gov/results/transients/ (Gehrels et al., 2004) in the 15–50 keV band during the 2023 outburst are shown in Fig. 1.
@@ 2.1 IXPE observations
The Imaging X-ray Polarimetry Explorer is a NASA mission in partnership with the Italian Space Agency (see a detailed description in Weisskopf et al., 2022), launched by a Falcon 9 rocket on 2021 December 9. There are three grazing incidence telescopes on board the observatory. Each telescope comprises an X-ray mirror assembly and a polarization-sensitive detector unit (DU) equipped with a gas-pixel detector (GPD; Soffitta et al., 2021; Baldini et al., 2021). These instruments provide imaging polarimetry in the 2–8 keV energy band with a time resolution better than 10 s.
IXPE observed J0243 three times during 2023 July 20–August 25 (ObsID 02250799), on July 20–22, August 10–11, and August 23–25, with the total exposures of 168, 77, and 131 ks per telescope, respectively (see Fig. 1). We refer to these observations as Obs. 1, 2, and 3, respectively. The data were processed with the ixpeobssim package version 31.0.1 (Baldini et al., 2022) using the CalDB version 20230702:v13. Source photons were collected in a circular region with a radius centered at the J0243 position. Because of the source brightness, the background was not subtracted following the recommendation by Di Marco et al. (2023). We performed the unweighted analysis (i.e., all events are taken into account independently of the quality of track reconstruction; Di Marco et al. 2022) of the IXPE data.
@@ 2.2 Timing
The event arrival times were corrected to the Solar System barycenter using the standard barycorr tool from the ftools666http://heasarc.gsfc.nasa.gov/ftools package (Blackburn, 1995) and accounting for the effects of binary motion using the orbital parameters from the Fermi/GBM777https://gammaray.nsstc.nasa.gov/gbm/science/pulsars.html (see Table 1). The spin period (of s) and the spin period derivative were then determined for each IXPE observation segment using an epoch-folding search method, and then refined using phase-connection technique, which allows the phase of each event within a given segment to be unambiguously defined. The results are presented in Table 4. However, the gaps between the individual observation segments coupled to the complex evolution of intrinsic spin frequency and pulse profiles together with the remaining uncertainties in orbital parameters preclude the direct connection of phases between the segments.
The absolute phase alignment between the segments was therefore determined independently. Inspection of the pulse profile shape observed by IXPE (see bottom panel of Fig. 9) reveals some common features: the minimum around phase zero, several subpeaks within the main peak, and a peak around phase 0.9 appearing late in the outburst. The observed phases of these features can thus be used to determine absolute phase offset for each segment such that all peaks and dips (determined through the fitting of Gaussian functions) appear at the same pulse phase. The residual scatter can then be used to assess the remaining uncertainty in the final phase alignment, which we estimate to be below 1%. The phase alignment can also be checked through the comparison of the observed IXPE pulse profiles with hard X-ray pulse profiles observed by Fermi/GBM.
To this end, we used the enhanced products provided by the GBM team containing Fourier coefficients of the pulse profiles for a set of time intervals with typical duration of 3 d. These are expected to change smoothly, and since Fermi data contains no data gaps, the individual pulse profiles can be aligned through the cross-correlation of subsequent time intervals. The resulting pulse profile evolution is presented in the top panel of Fig. 9. We see that changes similar to those revealed by IXPE also occur in the GBM data: the secondary peak around phase 0.9 appears to be present only at lower luminosities and disappears at higher luminosities (i.e., our initial alignment using IXPE data alone proves to be robust).
@@ 2.3 Spectral analysis
For the spectral analysis, Stokes spectra for Obs. 1, 2, and 3 were extracted using the xpbin tool’s PHA1 algorithm in ixpeobbsim, giving a set of three spectra (one for each DU) per observation. The energy spectra were binned to have at least 30 counts per energy channel. The spectra corresponding to the individual observations were fitted separately with the xspec package version 12.14.0 (Arnaud, 1996) using statistics. The uncertainties are given at the 68.3% confidence level. Considering the energy resolution and energy range of IXPE, we used a simplified model consisting of an absorbed power law plus a blackbody component to fit the J0243 spectra. To account for the interstellar absorption, a multiplicative component tbabs with the abundances from Wilms et al. (2000) was applied. A re-normalization constant const was introduced to account for possible discrepancies between effective areas of the different DUs and was fixed at unity for DU1. The full spectral model is thus
The results of the spectral fitting are shown in Fig. 3, and the best-fit parameters are given in Table 5. We see that the fit quality is not good, which is caused partially by systematic errors in the instrument response and partially by the presence of a weak iron line around 6.5 keV. However, this does not preclude us from measuring the flux in the IXPE band. For calculations of the total luminosity we estimated the bolometric correction factor for our IXPE observations from the NuSTAR spectra at similar luminosities during the 2017 outburst (Tsygankov et al., 2018; Bykov et al., 2022). The maximum bolometric luminosity of the 2023 outburst is erg s-1, a factor of 20 lower than the peak of the 2017 outburst (see penultimate row of Table 5).
We also studied variations in the spectrum over the pulse phase. Because the photon statistics do not allow us to use complex models, we fitted the phase-resolved spectra with a simple powerlaw model and fixed the hydrogen column density in the tbabs model at the best-fit phase-average values of 1.23, 1.01, and 1.56 cm-2 for Obs. 1, 2, and 3, respectively. Variations in the best-fit photon index with pulse phase are shown in Fig. 4b.
@@ 2.4 X-ray polarimetric analysis
The polarimetric parameters of J0243 were extracted utilizing the pcube algorithm (xpbin tool) in the ixpeobssim package (Baldini et al., 2022), which follows the description of Kislat et al. (2015). We derived the normalized Stokes parameters and , and subsequently the polarization degree PD= and polarization angle PA=. The uncertainties are given at the 68.3% confidence level unless stated otherwise.
The data were divided into 24 separate pulse-phase bins. The polarization was found to be low in the 2–3 keV band. This motivated us to use for the following analysis only the 3–8 keV data to reduce the noise. The pulse-phase dependence of the normalized Stokes parameters and is displayed in Fig. 4c and d, and their evolution on the -plane is shown in Fig. 5. The PD (shown in Fig. 4e) grows with time, from less than 8% in Obs. 1 to 12% in Obs. 2, and reaches a maximum of 24% in Obs. 3. The PA (Fig. 4f) has a double peak profile in all three observations. In Obs. 1 and 2 the amplitude of variations is about 100°,101010This is in strong disagreement with the recent analysis of Obs. 1 by Majumder et al. (2024), who used only five phase bins. while in Obs. 3 the data are consistent with two full revolution by 180° during the spin period. We do not see an obvious correlation between spectral shape and polarimetric properties.
@@ 2.5 Optical polarization studies
The IXPE observations were accompanied by optical polarimetric observations in the -band with the Robopol polarimeter located in the focal plane of the 1.3 m telescope of the Skinakas observatory (Greece). The observations were performed between 2023 July 28 and August 28 with multiple pointings. The data are presented in Table 11. We determined the intrinsic source polarization by subtracting the interstellar polarization using stars #1 and #2 (see Fig. 6), which are located 3′–4′ away from J0243 at a compatible distance of 5.7 kpc, as given by the Gaia EDR3 data (Bailer-Jones et al., 2021).
We also analyzed optical polarimetric measurements of the source during its previous outburst in 2017 obtained with the Double-Image Polarimeter (DIPol-2; Piirola et al., 2014) at the T60 telescope at Haleakala, Hawaii (see Table 14 in Appendix A). On average, 20 measurements of polarization were taken each night. We obtained the nightly averaged values using the clipping method (Kosenkov et al., 2017; Kosenkov, 2021). To estimate the interstellar polarization, we also observed field star #3 (see Fig. 6), which was reasonably bright and also has a parallax similar to that of the source.